### v1

# Metrics to Detect Small-Scale and Large-Scale Citation Orchestration

Abstract
Citation counts and related metrics have pervasive uses and misuses in academia and research appraisal, serving as scholarly influence and recognition measures. Hence, comprehending the citation patterns exhibited by authors is essential for assessing their research impact and contributions within their respective fields. Although the $h$-index, introduced by Hirsch in $2005$, has emerged as a popular bibliometric indicator, it fails to account for the intricate relationships between authors and their citation patterns. This limitation becomes particularly relevant in cases where citations are strategically employed to boost the perceived influence of certain individuals or groups, a phenomenon that we term “orchestration”. Orchestrated citations can introduce biases in citation rankings and therefore necessitate the identification of such patterns. Here, we use Scopus data to investigate orchestration of citations across all scientific disciplines. Orchestration could be small-scale, when the author him/herself and/or a small number of other authors use citations strategically to boost citation metrics like $h$-index; or large-scale, where extensive collaborations among many co-authors lead to high h-index for many/all of them. We propose three orchestration indicators: extremely low values in the ratio of citations over the square of the h-index (indicative of small-scale orchestration); extremely small number of authors who can explain at least $50\%$ of an author’s total citations (indicative of either small-scale or large-scale orchestration); and extremely large number of co-authors with more than $50$ co-authored papers (indicative of large-scale orchestration). The distributions, potential thresholds based on $1\%$ (and $5\%$) percentiles, and insights from these indicators are explored and put into perspective across science.
1 Introduction
Citation counts and related metrics are widely used and misused in scholarly evaluation [Bornmann, 2014]. They reflect the acknowledgment and utilization of research by the scientific community and play a significant role in assessing an author’s research impact and contributions within their respective fields [Costas et al., 2015, Ioannidis et al., 2014]. However, the conventional popular metrics used to quantify research impact may fall short in capturing the intricate relationships between authors and their citation patterns [Waltman and Van Eck, 2012].
For example, the $h$-index, introduced by Hirsch in 2005, has gained considerable popularity as a bibliometric indicator, measuring an author’s productivity and citation impact [Hirsch, 2005]. It represents the maximum number of an author’s papers that have received at least $h$ citations. While the $h$-index provides a straightforward measure of an author’s research impact, it overlooks the underlying complexities of citation patterns and the potential manipulation of citation metrics [Bornmann and Daniel, 2005]. Extensive criticism has been yielded against the $h$-index and its gaming potential [Kelly and Jennions, 2006, Chapman et al., 2019, Oravec, 2019].
In the scholarly landscape, there is an emerging concern regarding the strategic use of citations to enhance the influence of certain individuals or groups. This phenomenon, which can be referred to as “citation orchestration” occurs when authors deliberately include citations not solely to endorse relevant research but rather to boost the reputation or visibility of specific researchers or their affiliations [Didegah and Thelwall, 2013].
The $h$-index, for example, can be manipulated by self-citations placed strategically to boost it [Bartneck and Kokkelmans, 2011]. Self-citations may be the most superficial and readily detectable type of manipulation and some authors reach extremely high levels of self-citations, e.g. self-citations may account for more than $50\%$ or even $80\%$ of the citations that they receive [Van Noorden and Chawla, 2019, Ioannidis et al., 2020]. However, orchestration may also involve more subtle citation cartels: a small number of scientists may orchestrate to cite each other, regardless of whether they are co-authors or not [Fister et al., 2016, Perez et al., 2019]. Finally, with the advent of team science, it is becoming increasingly common for many teams to have a very large number of authors on each of the massive number of papers that they publish [Papatheodorou et al., 2008]. Such large-scale orchestration requires an in-depth analysis of the co-authorship patterns. All these behaviors and authorships and citation patterns are important to document so as to avoid a naive adoption of metrics like the $h$-index [Bornmann and Daniel, 2009, Ding et al., 2020].
Eventually, citation orchestration can distort citation rankings and compromise the fairness and accuracy of research evaluation processes [Waltman et al., 2011]. In order to understand the extent of citation orchestration and characterize its presence across science, here we use the entire Scopus database, which provides comprehensive coverage of scholarly literature across diverse disciplines and facilitates a multi-dimensional exploration of citation patterns [Burnham, 2006, Falagas et al., 2008]. We develop three readily accessible and interpretable metrics that may provide hints to possible citation orchestration behavior.
2 Methods
2.1 Proposed Metrics
The metrics are defined at the authors’ level. We propose three different metrics that may offer hints for the presence of inflated $h$-indices due to small-scale or large-scale citation orchestration:
$\mathbf{C/h^{2}}$:
The first metric is equal to the number of citations divided by the square of the author’s $h$-index. A small value of $C/h^{2}$ indicates that the citations of the author are coming from papers that explicitly increase the $h$-index. An extremely small value for this metric may suggest that the citations of the author may have been strategically placed in order to increase the $h$-index. This strategic placement may have been done either by the author himself/herself (self-citations) or by a small number of scientists who act as a coordinated citation cartel.
$\mathbf{A_{50\%C}}$:
This metric is equal to the number of citing authors who cumulatively explain $50\%$ of the citations of an author. An extremely small value of this metric may highlight either small-scale or large-scale orchestration. Small-scale orchestration results in a low $A_{50\%C}$ when a single author either cites himself/herself or is cited by some specific people within a citation cartel in order to promote his or her $h$-index. An extremely low $A_{50\%C}$ may be seen also with large-scale orchestration when a tightly-knit team effort dominates the scientific discipline and all/most authors are cited in all papers, a situation typical of nuclear and particle physics work affiliated with the European Organization of Nuclear Research. To calculate $A_{50\%C}$, we take into account the citing author who contributes the most citations to the examined author. Once this citing author $A_{1}$ is identified, we re-sort the remaining citing authors based on the remaining unexamined citing papers (i.e. excluding citing papers authored by $A_{1}$). The highest contributor $A_{2}$ is identified then and similarly removed. The process is iterated until citing author $A_{i}$, when at least $50\%$ of the citations are explained. This iterative process allows us to determine the number of citing authors who collectively explain at least $50\%$ of the examined authors’ citations while avoiding double-counting of citations from the same papers.
$\mathbf{A_{50}}$:
The metric is equal to the number of co-authors with shared co-authorship exceeding $50$ papers. If this number is large, this may highlight large-scale orchestration, where the authors belong to a big cluster of authors that heavily cite one another in work that they do jointly. Typically, this reflects the situation where prolific teams publish papers with massive co-authorship, and then these papers massively cite papers produced by the same team. This pattern is overwhelmingly seen in work done by the European Organization of Nuclear Research but this type of practice may also be appearing increasingly also in other scientific fields [Cronin, 2001].
2.2 Dataset
As part of the project, we were given access to the International Center for the Study of Research (ICSR) Lab [^1] https://www.elsevier.com/icsr/icsrlab, a cloud-based computational platform that allows its users to analyze large structured datasets such as the publication metadata of Scopus [^2] https://www.elsevier.com/solutions/scopus/how-scopus-works/metrics. The platform provides access to the publication metadata of $95,242,081$ papers in total as of $12^{th}$ of March 2023. These papers are authored by $49,255,352$ different author IDs in total. The precision and accuracy of the author ID files in Scopus have been well characterized [Kawashima and Tomizawa, 2015]; while some authors have their papers split into more than one ID file and some specific ID files contain papers by more than one author, the errors are very small to materially affect the distributions of the proposed metrics, and may be even less influential for the extreme tails that are likely to be most informative for orchestration.
2.3 Preprocessing
We only consider publications that are full papers, i.e. fall into the Scopus categories of articles, conference papers, or reviews, in concordance with previous work [Ioannidis et al., 2018], with a total of $82,694,786$ eligible papers. We next determine the field and subfield of study for each author using the Science-Metrix [^3] https://science-metrix.com/ [Archambault et al., 2011] classification. Each publication is assigned to one of 22 fields and one of 174 subfields. In case an author has published multiple papers and those papers belong to different fields, we choose the one where most of the author’s papers are assigned. In case of ties, we choose the field whose papers have received the largest number of citations. If there are still ties, we randomly select one of the fields. $2,889,779$ authors without specified subfield or field are excluded from all analyses. Table 1 shows the number of authors eligible for analysis in each field.
Field of study Number of authors Percentage Number of authors analyzed Percentage
Agriculture, Fisheries & Forestry $1,446,173$ $3.12\%$ $35,839$ $2.39\%$
Biology $1,586,467$ $3.42\%$ $64,737$ $4.33\%$
Biomedical Research $3,057,457$ $6.60\%$ $189,457$ $12.66\%$
Built Environment & Design $273,254$ $0.60\%$ $3,638$ $0.24\%$
Chemistry $2,464,113$ $5.31\%$ $91,721$ $6.13\%$
Clinical Medicine $15,594,844$ $33.63\%$ $579,366$ $38.71\%$
Communication & Textual Studies $328,254$ $0.71\%$ $1,617$ $0.11\%$
Earth & Environmental Sciences $1,345,726$ $2.90\%$ $58,067$ $3.88\%$
Economics & Business $909,251$ $2.00\%$ $20,776$ $1.39\%$
Enabling & Strategic Technologies $5,103,262$ $11.00\%$ $101,820$ $6.80\%$
Engineering $3,679,238$ $7.94\%$ $56,200$ $3.75\%$
General Arts, Humanities & Social Sciences $5,992$ $0.01\%$ $2$ $0.00\%$
General Science & Technology $30,121$ $0.06\%$ $77$ $0.01\%$
Historical Studies $241,666$ $0.52\%$ $1,550$ $0.10\%$
Information & Communication Technologies $4,094,210$ $8.83\%$ $57,163$ $3.82\%$
Mathematics & Statistics $410,548$ $0.89\%$ $9,605$ $0.64\%$
Philosophy & Theology $111,270$ $0.24\%$ $433$ $0.03\%$
Physics & Astronomy $2,649,145$ $5.71\%$ $174,028$ $11.63\%$
Psychology & Cognitive Sciences $487,914$ $1.10\%$ $18,264$ $1.22\%$
Public Health & Health Services $1,221,920$ $2.64\%$ $19,928$ $1.33\%$
Social Sciences $1,300,068$ $2.80\%$ $12,372$ $0.83\%$
Visual & Performing Arts $25,800$ $0.06\%$ $20$ $0.00\%$
Table 1: Number and percentage of total Scopus author IDs and of Scopus author IDs eligible for analysis (those with at least 5 full papers and at least 1000 citations).
We focus on a subset of authors who have made a significant contribution to the research community and satisfy the following two conditions: (1) they have published more than $5$ full papers, and (2) their papers have received at least $1,000$ citations in total. Eventually, $1,496,680$ authors satisfy the above conditions. Their numbers per scientific field are also shown in Table 1.
2.4 Analyses of distributions for proposed indicators
For each of the three proposed indicators, we generate the distribution of values for the $1,496,680$ eligible authors. We focus on the left-side tails (very small values) for the first two indicators and the right-side tails for the third one. There is no absolute cut-off that can separate authors where orchestration has been done on purpose. The most extreme values are more likely to reflect orchestration, but orchestration may have happened to a various extend also for citation profiles of authors that have less extreme values. Here we focus primarily on the lowest $1\%$ percentile for $C/h^{2}$ and $A_{50\%C}$ and the highest $1\%$ for $A_{50}$ metric. We also present values for the $5\%$ threshold as well as median and interquartile range values for each of the three indicators. We also show how extreme $1\%$ percentile authors are distributed across the main fields of science, so as to demonstrate whether these patterns are enriched in specific fields. For $A_{50\%C}$ and $A_{50}$ specifically, we provide analyses that exclude the authors categorized in the field of Physics and Astronomy, since their tails of interest are otherwise overwhelmed by nuclear and particle physics authors in work affiliated with the European Organization for Nuclear Research.
3 Results
3.1 Distributions of indicators
The distribution of the three proposed indicators appears in Figure 1 and Figure 2 zooms out specifically on the lower $5\%$ percentile of the first two indicators and the upper $5\%$ percentile of the third orchestration indicator. The distributions are right-skewed. For the $A_{50\%C}$ and $A_{50}$ metrics, we provide also distributions excluding the field of Physics and Astronomy so as to allow revealing better the patterns for all other fields.
3.2 Metric $C/h^{2}$
For the metric $C/h^{2}$, the $1^{st}$ percentile value is $2.45$ and the $5^{th}$ percentile value is $2.76$. Comparatively, the median is $4.11$ and the interquartile range is $3.36$ to $6.11$. Here, we focus on the 1% percentile ($14,967$ authors) that may have the strongest hint for small-scale orchestration. Table 2 shows that the scientific field allocation of these $14,967$ authors is largely similar to the allocation of all examined authors, with some exceptions. Specifically, there are proportionally more than $1.5$-fold higher representation of authors with extremely low $C/h^{2}$ values compared to all examined authors in 3 fields: Agriculture, Fisheries and Forestry, Chemistry, and Earth and Environmental Sciences.
(a) The distribution of the metric $C/h^{2}$ (authors with values above 20 are not shown)
(b) The distribution of the metric $A_{50\%C}$ (authors with values above 200 are not shown)
(c) The distribution of the metric $A_{50}$ (authors with values of 0 and above 40 are not shown).
(d) The distribution of the metric $A_{50\%C}$ (excluded authors with values exceeding 200 and those in the field of Physics and Astronomy).
(e) The distribution of the metric $A_{50}$ (excluded authors with values above 40 and those in the field of Physics and Astronomy).
Figure 1: The distribution of all the proposed metrics
(a) The distribution of the metric $C/h^{2}$ on the lowest 5%
(b) The distribution of the metric $A_{50\%C}$ on the lowest 5%
(c) The distribution of the metric $A_{50}$ on the upper 5% (authors with values above 100 are not shown)
Figure 2: The distribution of all the proposed metrics on the lowest or upper 5%
Field of study Percentage of examined authors Percentage of lowest 1% authors of $C/h^{2}$ metric Percentage of lowest 1% authors of metric $A_{50\%C}$ (excluding Physics & Astronomy) Percentage of upper 1% authors of metric $A_{50}$ (excluding Physics & Astronomy)
Agriculture, Fisheries & Forestry $2.39\%$ $5.74\%$ $2.19\%$ $1.43\%$
Biology $4.33\%$ $5.99\%$ $4.34\%$ $0.75\%$
Biomedical Research $12.66\%$ $12.19\%$ $4.76\%$ $9.03\%$
Built Environment & Design $0.24\%$ $0.25\%$ $0.44\%$ $0.02\%$
Chemistry $6.13\%$ $13.80\%$ $15.47\%$ $3.07\%$
Clinical Medicine $38.71\%$ $29.20\%$ $17.09\%$ $73.29\%$
Communication & Textual Studies $0.11\%$ $0.03\%$ $0.26\%$ $0.00\%$
Earth & Environmental Sciences $3.88\%$ $5.93\%$ $4.82\%$ $1.34\%$
Economics & Business $1.39\%$ $0.62\%$ $2.13\%$ $0.00\%$
Enabling & Strategic Technologies $6.80\%$ $7.96\%$ $10.84\%$ $7.31\%$
Engineering $3.75\%$ $4.64\%$ $12.18\%$ $1.17\%$
General Arts, Humanities & Social Sciences $0.00\%$ $0.00\%$ $0.00\%$ $0.00\%$
General Science & Technology $0.01\%$ $0.00\%$ $0.00\%$ $0.00\%$
Historical Studies $0.10\%$ $0.08\%$ $0.24\%$ $0.00\%$
Information & Communication Technologies $3.82\%$ $1.06\%$ $11.00\%$ $1.59\%$
Mathematics & Statistics $0.64\%$ $0.48\%$ $9.44\%$ $0.04\%$
Philosophy & Theology $0.03\%$ $0.00\%$ $0.16\%$ $0.00\%$
Physics & Astronomy $11.63\%$ $9.77\%$- --
Psychology & Cognitive Sciences $1.22\%$ $0.84\%$ $1.99\%$ $0.25\%$
Public Health & Health Services $1.33\%$ $0.92\%$ $0.96\%$ $0.67\%$
Social Sciences $0.83\%$ $0.46\%$ $1.68\%$ $0.03\%$
Visual & Performing Arts $0.00\%$ $0.00\%$ $0.02\%$ $0.00\%$
Table 2: Authors with extreme metrics: Distribution across the 22 main fields of science.
(a) Distribution of the citations, considering only whoever author has an $C/h^{2}$ value less than $2.45$ (authors with citations above $5,000$ are not shown).
(b) The distribution of the number of papers published by authors with an $C/h^{2}$ value on the lowest 1% (authors with more than 200 papers are not shown).
(c) Distribution of the $h$-index of the authors that have an $C/h^{2}$ value less than $2.45$ (authors with $h$-index above 60 are not shown).
Figure 3: The distribution of citations, number of papers and $h$-index of the authors that belong to the lowest 1% of metric $C/h^{2}$.
Figure 3 depicts the distribution of citations, the number of published papers, and the $h$-index, respectively, for this group of $14,717$ authors with the lowest $C/h^{2}$ metric. These are highly productive authors (median $63$ papers, the interquartile range is $50$ to $81$ publications) and they all have $h$-indices exceeding $21$, while some of them have even very large $h$-indices.
3.3 Metric $A_{50\%C}$
For the $A_{50\%C}$ metric, we observed that $73\%$ ($4,339$ out of $5,881$) of authors with the lowest $1\%$ percentile values of $A_{50\%C}$ belong to Physics & Astronomy (of those, $3,808$ belong to the subfield of nuclear and particle physics). Therefore, we excluded Physics & Astronomy authors in order to examine better the distribution for the rest of science. All results that follow retain analyses excluding Physics & Astronomy. For the $A_{50\%C}$ metric, the $1^{st}$ percentile is $5$ while the median is $36$, and the interquartile range is $21$ to $60$. The $11,277$ authors who belong to this lowest $1\%$ percentile (metric value less than $5$) are more likely to engage in self-citation or be part of networks that highly cite each other. Table 2 shows the disciplines to which these $11,277$ authors belong. There is a more than 1.5-fold higher representation of authors with exceptionally low $A_{50\%C}$ values compared to all examined authors in $10$ specific fields: Chemistry, Enabling & Strategic Technologies, Engineering, Historical Studies, Information & Communication Technologies, Mathematics & Statistics, Philosophy and Theology, Psychology and Cognitive Sciences, Social Sciences, and Visual and Performing Arts.
Figure 4 depicts the distribution of citations, the number of published papers, and the h-index, respectively, for this group of $11,277$ authors with the lowest $1\%$ percentile for the $A_{50\%C}$ metric. These are mostly very productive authors (median $104$ papers, interquartile range $71$ to $153$ publications) and they mostly ($59\%$ of these authors) have $h$-indices exceeding $20$, while some of them have even large $h$-indices.
(a) Distribution of the citations, considering only whoever author has an $A_{50\%C}$ value less than $5$ (authors with citations above $10,000$ are not shown).
(b) The distribution of the number of papers published by authors with an $A_{50\%C}$ value on the lowest 1% (authors with more than $500$ papers are not shown).
(c) Distribution of the $h$-index of the authors that have an $A_{50\%C}$ value less than $5$ (authors with $h$-index above $60$ are not shown).
Figure 4: The distribution of citations, number of papers and $h$-index of the authors that belong to the lowest 1% of metric $A_{50\%C}$.
3.4 Metric $A_{50}$
Metric $A_{50}$ aims to identify authors that heavily collaborate with each other, essentially functioning as a metric for detecting large-scale orchestration. Among the $1,496,680$ considered authors, we exclude the authors that belong to the field of ”Physics & Astronomy”, as it is widely recognized for its consistent involvement in large-scale collaborative endeavors, leading to a potential bias in the analysis. In particular, we observed that around $99\%$ of authors with the largest $1\%$ percentile values of $A_{50}$ belong to Physics & Astronomy. By excluding those authors from our analysis, we aim to mitigate this bias and gain a clearer understanding of metric $A_{50}$ in a broader context. Thus, we end up examining $1,322,652$ authors. Among them, $242,281$ have a value greater than zero for the $A_{50}$ metric, they have at least $1$ co-author with whom they have co-authored more than $50$ papers.
For this metric the $99^{th}$ percentile value is $7$, and the $95^{th}$ percentile is $2$. The median is equal to 00 as well as the interquartile range. Here, we focus on the top $1\%$ of authors who possess an $A_{50}$ value exceeding $7$, as we place particular emphasis on exceptional cases where an author has collaborated with a diverse set of individuals more than $50$ times, providing the strongest indication for large-scale orchestration. Table 2 indicates the allocation of these $12,015$ authors across scientific disciplines. A strong predilection is shown for authors within Clinical Medicine, as they represent $73.29\%$ of these $12,015$ authors, as opposed to only $38.71\%$ among all authors. No other field has an enrichment among the $12,015$ authors. Supplementary table 4 shows the breakdown of Clinical Medicine authors into subfields. As shown, allergy, arthritis & rheumatism, cardiovascular system & hematology, gastroenterology & hepatology, general clinical medicine, geriatrics, and psychiatry subfields are 1.5-fold or more enriched among those with $A_{50}$ values exceeding the highest $1\%$ percentile.
Figure 5 depicts the distribution of citations, the number of published papers, and the $h$-index, respectively, for this group of $12,015$ authors with the largest $A_{50}$ metric. These are highly productive authors (median $596$ papers, interquartile range $315$ to $1061$) and their $h$-indices are typically very high or even extraordinarily high (median $56$, interquartile range $38$ to $80$, $12.8\%$ of these authors with $h$-index¿100).
(a) Distribution of the citations, considering only whoever author has an $A_{50}$ value greater than $7$ (authors with citations above $100,000$ are not shown).
(b) The distribution of the number of papers published by authors with an $A_{50}$ value on the upper 1% (authors with more than $2,000$ papers are not shown).
(c) Distribution of the $h$-index of the authors that have an $A_{50}$ value greater than $7$ (authors with $h$-index above $200$ are not shown).
Figure 5: The distribution of citations, number of papers and $h$-index of the authors that belong to the upper 1% of metric $A_{50}$.
3.5 Co-existence of extreme values for the metrics
We examined the co-existence of extreme values for the three metrics (lowest $1\%$ percentile for the first two, highest $1\%$ percentile for the third one) among all $1,322,652$ authors excluding Physics and Astronomy. As shown in Table 3, there was strong co-existence for extremely low values of $C/h^{2}$ and $A_{50\%C}$ (odds ratio $6.4$, $95\%$ confidence interval $5.9-6.9$), modest coexistence for extremely low values of $A_{50\%C}$ and extremely high values of $A_{50}$ (odds ratio $1.5$, 95% confidence interval $1.3-1.7$), and extremely low values of $C/h^{2}$ very rarely co-occurred with extremely high values of $A_{50}$ (odds ratio $0.09$, $95\%$ confidence interval $0.05-0.16$).
Lowest percentile of $A_{50\%C}$ Highest percentile of $A_{50}$
YES NO YES NO
Lowest percentile of $C/h^{2}$ YES $659$ $12,566$ $11$ $13,214$
NO $10,618$ $1,298,809$ $12,004$ $1,297,423$
Lowest percentile of $A_{50\%C}$ YES $151$ $11,126$
NO $11,864$ $1,299,511$
Table 3: Co-existence of small- and large-scale orchestration indicators
4 Discussion
The phenomenon of orchestrated citations may markedly enhance the influence of specific individuals or groups. Using science-wide data from Scopus, we present three readily accessible metrics that may offer hints to orchestration patterns. This work may contribute to the ongoing efforts to enhance the assessment of research impact and recognize the contributions of authors within their respective fields [Ioannidis and Maniadis, 2023b, Ioannidis and Maniadis, 2023a]. We found that authors with hints of orchestration based on these 3 metrics have enriched presence in specific scientific fields. Importantly, these authors almost always had high or even extremely high h-indices. These high $h$-indices would be very misleading and would need to be interpreted in the light of the authorship and citation patterns to place them in a more proper context.
Authors at the extreme tails of these metrics may or may not engage in individual or group practices that lead to inflated authorship contributions and inflated perceived impact. Moreover, most authors who do occasionally engage in practices that lead to inflated authorship contributions [Masic and Jankovic, 2021] or inflated perceived impact would not have such extreme values in these metrics. We propose to use these metrics with a very stringent threshold, such as 1% percentile, where chance is unlikely to generate such extremes in the absence of an organized plan or group arrangement regarding the allocation of authorship and/or citations.
However, any claims of ethical wrong doing should be withheld in the absence of additional evidence. Each author with hints of extreme orchestration would need to be examined carefully as to the nature of their collaborations, teamwork arrangements and standards in their environment, and citation practices in the networks of scientists who cite them. These metrics should not be seen as synonymous with unethical practices but as hints for further appraisal. Many of these scientists may have perfectly legitimate standards for their field, e.g. massive co-authorship in their teamwork may be the accepted norm, as in the case of nuclear and particle physics work. Nevertheless, even in these cases, the extreme nature of orchestration should be noted and the citation metrics of these authors should be properly adjusted (e.g. for co-authorship [Ioannidis, 2008] and interpreted. This is essential in order to avoid unfair comparisons against other scientists in the same or neighboring fields who have different work and publication arrangements. Otherwise, comparisons would be grossly unfair.
We observed that the two metrics of small-scale orchestration tended to coexist far more common than chance. Authors with double indicators for small-scale orchestration should be even more carefully scrutinized. The presence of unjustified extreme self-citations or citation cartels should be assessed in these cases. This requires an in-depth assessment of the nature, relevance, and justification of citations. Future work may also explore whether such behavior is linked to other unethical or spurious publication practices, such as paper mills [Christopher, 2021], hijacked journals [Dadkhah and Borchardt, 2016, Abalkina, 2021], hijacked citations [Dadkhah, 2016] and other subversive arrangements.
The two indicators of large-scale orchestration also tended to co-exist more often than chance. In authors with large-scale orchestration, sometimes astonishingly high citation counts and $h$-indices can be reached, but they are largely meaningless. These authors should be assessed for the nature of their contribution to the massively collaborative teamwork where they participate. Several unethical behaviors may co-exist in some cases of large-scale orchestration, e.g. scientists may be placed as authors with gift authorship [Smith, 1994, Teixeira da Silva and Dobránszki, 2016]. This is a particular problem with people who acquire power as chairs/heads and exhibit a massive acceleration of their research productivity (at a time when administrative duties should probably have cut their productivity). Moreover, while teamwork is a good research practice, in principle, inflation of authorship within teams may signal salami-slicing, and a poor sense of accountability and may reflect badly on the entire team/consortium.
Some limitations of our work should be discussed. First, we only focused on authors with at least 1000 citations. Orchestration may also affect scientists who have fewer citations. Scientists with fewer than 1000 citations are the vast majority of the scientific workforce. However, their influence in the scientific literature is limited compared with the more influential sample that we examined. Moreover, it is common for self-citations to be high in the early career of scientists when they have limited citations and they try to establish themselves with their work [Seeber et al., 2019]. This would possibly frequently raise signals of small-scale orchestration in early career authors that have a very different meaning compared with established authors with many citations. Second, inaccuracies in Scopus may affect the calculation of these metrics for these scientists. Inaccuracies are overall low across Scopus, but in the case of an in-depth assessment of the behavior of specific authors suspected of unethical orchestration practices, it is important to verify at the first step that their publication and citation profile is accurate, so as to avoid mischaracterizations. Third, we did not employ here more complex methods of community detection, centrality analysis, and network visualization [Jolly et al., 2020, Fister Jr et al., 2016], but these methods may also be utilized when unethical orchestration is suspected and in-depth evaluation of specific author networks is needed.
Overall our work reveals that many authors with high or even exceptionally high traditional citation metrics such as the h-index have strong hints of orchestration practices. This distortion should be taken into account in using citation metrics for evaluation and reward purposes.
Appendix A Supplementary material
Subfield Number of eligible authors Percentage of eligible authors Number of authors that belong to highest 1% percentile Percentage over the authors of highest 1% percentile
allergy $2,922$ $0.50\%$ $93$ $1.06\%$
anesthesiology $4,562$ $0.79\%$ $16$ $0.18\%$
arthritis & rheumatology $6,614$ $1.14\%$ $168$ $1.91\%$
cardiovascular system & hematology $61,481$ $10.61\%$ $1,462$ $16.60\%$
dentistry $8,488$ $1.47\%$ $22$ $0.25\%$
dermatology & venereal diseases $7,218$ $1.25\%$ $77$ $0.87\%$
emergency & critical care medicine $4,602$ $0.79\%$ $35$ $0.39\%$
endocrinology & metabolism $25,727$ $4.44\%$ $373$ $4.24\%$
environmental & occupational health $1,289$ $0.22\%$ $9$ $0.10\%$
gastroenterology & hepatology $18,708$ $3.23\%$ $500$ $5.68\%$
general & internal medicine $43,032$ $7.43\%$ $448$ $5.09\%$
general clinical medicine $287$ $0.05\%$ $9$ $0.10\%$
geriatrics $937$ $0.16\%$ $23$ $0.26\%$
immunology $46,471$ $8.02\%$ $470$ $5.34\%$
legal & forensic medicine $601$ $0.10\%$ $5$ $0.06\%$
neurology & neurosurgery $92,605$ $15.98\%$ $1,096$ $12.45\%$
nuclear medicine & medical imaging $15,719$ $2.71\%$ $143$ $1.62\%$
obstetrics & reproductive medicine $14,860$ $2.56\%$ $139$ $1.58\%$
oncology & carcinogenesis $95,554$ $16.49\%$ $2,041$ $23.18\%$
ophthalmology & optometry $10,206$ $1.76\%$ $65$ $0.74\%$
orthopedics $12,512$ $2.16\%$ $193$ $2.19\%$
otorhinolaryngology $4,199$ $0.72\%$ $19$ $0.22\%$
pathology $1,130$ $0.19\%$ $4$ $0.05\%$
pediatrics $9,012$ $1.56\%$ $56$ $0.64\%$
pharmacology & pharmacy $21,038$ $3.63\%$ $64$ $0.73\%$
psychiatry $17,236$ $2.97\%$ $481$ $5.46\%$
respiratory system $11,667$ $2.01\%$ $179$ $2.03\%$
sport sciences $4,521$ $0.78\%$ $25$ $0.28\%$
surgery $18,604$ $3.21\%$ $303$ $3.44\%$
tropical medicine $3,024$ $0.52\%$ $25$ $0.28\%$
urology & nephrology $14,419$ $2.49\%$ $263$ $2.99\%$
Table 4: The break-down of Clinical Medicine authors, who belong to the highest 1% of metric $A_{50}$, into subfields